Shift report 20000714 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Fri Jul 14 2000 - 07:48:14 EDT

  • Next message: Konstantin Olchanski: "Re: Shift report 20000714 00:00-08:00"

    ------------------------------------------------------------
    Date: Fri Jul 14 07:48:14 2000, Shift: 00:00-08:00
    Supervisor: Michael Murray, with: all alone with only the deer to talk to
    Log book#2 pages 151 - 155
    
    Shift Summary:
    ------------------------------------------------------------
    The accelerator group practised ramping the beam to 
    transition troughout the night. The energy was never 
    above 22GeV. The beam was lost, or dumped, ever 20 minutes
    or so.
    
    Generally the experiment seems in good shape after the
    break. ZDC, BB, Mult, H1, H2 and TPCs looked OK. 
    ZDC ADC has 2 channel empty. They were moved to another
    ADC for a test. The beam - beam counter has 2 empty TDC
    spectra for Right Outer Tubes, 17 and 19. The
    TOFW lost the TDC in slot 10 and this confused the DAQ
    so we got no TOFW data through the night. For the 
    Si every 8th channel, 7, 15, 23 ... 191 showed only 
    a pedestal. Most channels looked good with narrow pedestals
    at about 20 counts and maybe 1 mip peaks. However channels
    80 to 87 and 112 to 190 have ADC spectra that start at
    200 counts.
    
    It would be useful if their was a quick way to scan the 
    Si ADCs in the monitoring. Also the TOFW histogram titles
    are not readable.
    
    7.45am Beam now at 65GeV. First time all night. 
    Runs 1764-1767  can be used to debug detectors but do 
    not contain any useful physics data.
     
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Fri Jul 14 2000 - 08:26:20 EDT